\begin{thebibliography}{10}

\bibitem{JeMoSo1999}
U.~D. Jentschura, P.~J. Mohr, and G. Soff, Phys. Rev. Lett. {\bf 82},  53
  (1999).

\bibitem{JeMoSo2001pra}
U.~D. Jentschura, P.~J. Mohr, and G. Soff, Phys. Rev. A {\bf 64},  042512
  (2001).

\bibitem{SaYe1990}
J. Sapirstein and D.~R. Yennie,  in {\em Quantum Electrodynamics}, edited by T.
  Kinoshita (World Scientific, Singapore, 1990), pp.\ 560--672.

\bibitem{Be1947}
H.~A. Bethe, Phys. Rev. {\bf 72},  339  (1947).

\bibitem{Fe1948}
R.~P. Feynman, Phys. Rev. {\bf 74},  1430  (1948).

\bibitem{Fe1949}
R.~P. Feynman, Phys. Rev. {\bf 76},  769  (1949).

\bibitem{FrWe1949}
J.~B. French and V.~F. Weisskopf, Phys. Rev. {\bf 75},  1240  (1949).

\bibitem{KrLa1949}
N.~M. Kroll and W.~E. Lamb, Phys. Rev. {\bf 75},  388  (1949).

\bibitem{Sc1949}
J. Schwinger, Phys. Rev. {\bf 75},  898  (1949).

\bibitem{FuMiTo1949}
H. Fukuda, Y. Miyamoto, and S. Tomonaga, Prog. Theor. Phys. (Kyoto) {\bf 4},
  47  (1949).

\bibitem{Ba1951}
M. Baranger, Phys. Rev. {\bf 84},  866  (1951).

\bibitem{KaKlSc1952}
R. Karplus, A. Klein, and J. Schwinger, Phys. Rev. {\bf 86},  288  (1952).

\bibitem{BaBeFe1953}
M. Baranger, H.~A. Bethe, and R.~P. Feynman, Phys. Rev. {\bf 92},  482  (1953).

\bibitem{FrYe1958}
H.~M. Fried and D.~R. Yennie, Phys. Rev. {\bf 112},  1391  (1958).

\bibitem{FrYe1960}
H.~M. Fried and D.~R. Yennie, Phys. Rev. Lett. {\bf 4},  583  (1960).

\bibitem{La1960}
A.~J. Layzer, Phys. Rev. Lett. {\bf 4},  580  (1960).

\bibitem{La1961a}
A.~J. Layzer, J. Math. Phys. {\bf 2},  292  (1961).

\bibitem{La1961b}
A.~J. Layzer, J. Math. Phys. {\bf 2},  308  (1961).

\bibitem{ErYe1965a}
G.~W. Erickson and D.~R. Yennie, Ann. Phys. (N.Y.) {\bf 35},  271  (1965).

\bibitem{ErYe1965b}
G.~W. Erickson and D.~R. Yennie, Ann. Phys. (N.Y.) {\bf 35},  447  (1965).

\bibitem{Er1971}
G.~W. Erickson, Phys. Rev. Lett. {\bf 27},  780  (1971).

\bibitem{Sa1981}
J. Sapirstein, Phys. Rev. Lett. {\bf 47},  1723  (1981).

\bibitem{Pa1993}
K. Pachucki, Ann. Phys. (N.Y.) {\bf 226},  1  (1993).

\bibitem{KlMa1973}
S. Klarsfeld and A. Maquet, Phys. Lett. B {\bf 43},  201  (1973).

\bibitem{BeBrSt1950}
H.~A. Bethe, L.~M. Brown, and J.~R. Stehn, Phys. Rev. {\bf 77},  370  (1950).

\bibitem{Ha1956}
J.~M. Harriman, Phys. Rev. {\bf 101},  594  (1956).

\bibitem{ScTi1959}
C. Schwartz and J.~J. Tieman, Ann. Phys. (N.Y.) {\bf 6},  178  (1959).

\bibitem{Li1968}
M. Lieber, Phys. Rev. {\bf 174},  2037  (1968).

\bibitem{Hu1969}
R.~W. Huff, Phys. Rev. {\bf 186},  1367  (1969).

\bibitem{GoDr2000}
S.~P. Goldman and G.~W.~F. Drake, Phys. Rev. A {\bf 61},  052513  (2000).

\bibitem{MoTa2000}
P.~J. Mohr and B.~N. Taylor, Rev. Mod. Phys. {\bf 72},  351  (2000).

\bibitem{MoTaPriv2004}
P.~J. Mohr and B.~N. Taylor, CODATA recommended values of the fundamental
  physical constants: 2002 (to be published).

\bibitem{MoKi1992}
P.~J. Mohr and Y.~K. Kim, Phys. Rev. A {\bf 45},  2727  (1992).

\end{thebibliography}

\documentclass[twocolumn,aps,prd,showpacs,superscriptaddress,floatfix]{revtex4}

\usepackage{bm}
\usepackage{dcolumn}
\usepackage{graphicx}
\usepackage{epsf}
\usepackage{epsfig}

\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}

\newcommand{\addrFR}{Theoretische Quantendynamik,
Physikalisches Institut der Universit\"{a}t Freiburg,\\
Hermann--Herder--Stra\ss{}e 3, 79104 Freiburg im Breisgau, Germany}

\newcommand{\addrNIST}{National Institute of Standards and Technology,
Gaithersburg, Maryland 20899--8401}

\begin{document}

\bibliographystyle{myprsty}

\title{Electron Self Energy for Higher Excited $S$ Levels}

\author{Ulrich~D.~Jentschura}
\affiliation{\addrFR}
\affiliation{\addrNIST}

\author{Peter~J.~Mohr}
\affiliation{\addrNIST}

\begin{abstract}
A nonperturbative numerical evaluation of the one-photon electron self
energy for the $3S$ and $4S$ states with 
charge numbers $Z=1$ to 5 is described. 
The numerical results are in agreement with known terms in 
the expansion of the self energy in
powers of $Z\alpha$.
\end{abstract}

\pacs{12.20.Ds, 31.30.Jv, 06.20.Jr, 31.15.-p}

\maketitle

In this brief report, we consider the one-loop self-energy shift 
which is the dominant radiative correction to the 
energy of hydrogenic bound states.
For high-precision spectroscopy, $S$ states are rather
important because they can be excited from the ground state
via Doppler-free two-photon spectroscopy.
We calculate the self-energy numerically to high accuracy
for $3S$ and $4S$ states (nuclear charge number
$Z=1,\dots,5$). We follow the approach previously 
outlined for $1S$~(Ref.~\cite{JeMoSo1999}) and 
$2S$ and $2P$ states~(Ref.~\cite{JeMoSo2001pra}).

The natural
unit system with $\hbar = c = m_{\rm e} = 1$ and $e^2 = 4\pi\alpha$ is
employed, as is customary in bound-state quantum
electrodynamics.
The (real part of the) energy shift $\Delta E_{\rm SE}$ due to the
electron self-energy radiative correction is usually written 
as~\cite{SaYe1990}
\begin{equation}
\label{ESEasF}
\Delta E_{\rm SE} = \frac{\alpha}{\pi} \, 
\frac{(Z \alpha)^4 \, m_{\rm e}}{n^3} \, 
F(nl_j,Z\alpha) \,,
\end{equation}
where $F$ is a dimensionless quantity. 
In writing the expression $F(nl_j,Z\alpha)$, we follow the usual
spectroscopic notation for an atomic state with 
principal quantum number
$n$, orbital angular momentum $l$ and 
total electron angular momentum $j$.  

\begin{table}[htb]
\begin{center}
\begin{minipage}{7.7cm}
\begin{center}
\caption{\label{tableF3S} Numerical results for
the scaled self-energy function
$F$ ($3S$ state) and the self-energy remainder function $G_{\rm SE}$,
in the regime of low nuclear charge numbers $Z$.}
\begin{tabular}{l@{\hspace*{0.5cm}}r@{.}l@{\hspace*{0.5cm}}r@{.}l}
\hline
\hline
$Z$ &
\multicolumn{2}{l}{\rule[-3mm]{0mm}{8mm}
  $F(3{S}_{1/2},Z\alpha)$} &
\multicolumn{2}{l}{\rule[-3mm]{0mm}{8mm}
  $G_{\rm SE}(3{S}_{1/2},Z\alpha)$}\\
\hline
\rule[-1mm]{0mm}{6mm}
$1$ & $10$ & $605~614~22(5)$ & $-31$ & $047~7(9)$ \\
\rule[-1mm]{0mm}{6mm}
$2$ &  $8$ & $817~615~14(5)$ & $-30$ & $512~6(2)$ \\
\rule[-1mm]{0mm}{6mm}
$3$ &  $7$ & $794~461~17(5)$ & $-30$ & $022~7(1)$ \\
\rule[-1mm]{0mm}{6mm}
$4$ &  $7$ & $083~612~42(5)$ & $-29$ & $564~53(6)$ \\
\rule[-3mm]{0mm}{8mm}
$5$ &  $6$ & $543~385~98(5)$ & $-29$ & $130~61(4)$ \\
\hline
\hline
\end{tabular}
\end{center}
\end{minipage}
\end{center}
\end{table}

The leading terms 
in the semi-analytic expansion of $F(n{S}_{1/2},Z\alpha)$
about $Z\alpha=0$ read
\begin{eqnarray}
\label{defFLOnS}
\lefteqn{F(n{S}_{1/2},Z\alpha) = 
A_{41}(n{S}_{1/2}) \, \ln(Z \alpha)^{-2}} \nonumber\\[2ex]
& & \;\; + A_{40}(n{S}_{1/2}) +
(Z \alpha) \, A_{50}(n{S}_{1/2}) \nonumber\\[2ex]
& & \;\; + \, 
(Z \alpha)^2 \, 
\left[A_{62}(n{S}_{1/2}) \, \ln^2(Z \alpha)^{-2} \right. 
\nonumber\\[2ex]
& & \;\;\;\; \left. + A_{61}(n{S}_{1/2}) \,\ln(Z \alpha)^{-2} + 
G_{\rm SE}(n{S}_{1/2},Z\alpha) \right]\,.
\end{eqnarray}
The $A$ coefficients have two indices, the first of which denotes the
power of $Z\alpha$ [including those powers implicitly contained in
Eq.~(\ref{ESEasF})], while the second index denotes the power of the
logarithm $\ln(Z \alpha)^{-2}$.  

We now list the analytic coefficients and the
Bethe logarithms relevant to the atomic states under
investigation~\cite{Be1947,Fe1948,Fe1949,FrWe1949,KrLa1949,Sc1949,FuMiTo1949,%,
Ba1951,KaKlSc1952,BaBeFe1953,FrYe1958,FrYe1960,La1960,La1961a,La1961b,%,
ErYe1965a,ErYe1965b,Er1971,Sa1981,Pa1993}, 
\begin{subequations}
\begin{eqnarray}
\label{coeffs1S12}
A_{41}(n{S}_{1/2}) & = & \frac{4}{3}\,, \\[1ex]
A_{40}(n{S}_{1/2}) & = & 
  \frac{10}{9} - \frac{4}{3} \, \ln k_0(n{S})\,, 
\\[1ex]
A_{50}(n{S}_{1/2}) & = & 
4\pi\,\left[\frac{139}{128} - \frac{1}{2}\,\ln 2\right]\,,
\\[1ex]
A_{62}(n{S}_{1/2}) & = & -1\,.
\end{eqnarray}
\end{subequations}
$A_{61}$-coefficients read
\begin{subequations}
\begin{eqnarray}
A_{61}(1{S}_{1/2}) & = & \frac{21}{20} + \frac{28}{3} \, \ln 2 \,,\\[1ex]
A_{61}(2{S}_{1/2}) & = & \frac{67}{30} + \frac{16}{3} \, \ln 2 \,,\\[1ex]
A_{61}(3{S}_{1/2}) & = & \frac{6163}{1620} + \frac{28}{3} \, \ln 2 - 4\,\ln 3\,,
\\[1ex]
A_{61}(4{S}_{1/2}) & = & \frac{4}{3} \, \ln 2 + \frac{391}{80}\,.
\end{eqnarray}
\end{subequations}
The Bethe logarithms $\ln k_0(n{S})$ are 
known~\cite{KlMa1973,BeBrSt1950,Ha1956,ScTi1959,Li1968,Hu1969},
and we here present a re-evaluation, 
\begin{subequations}
\begin{eqnarray}
\label{BetheLog1S}
\ln k_0(1{S}) &=& 2.984~128~555~765~497~611(1)\,,\\[1ex]
\label{BetheLog2S}
\ln k_0(2{S}) &=& 2.811~769~893~120~563~520(1)\,,\\[1ex]
\label{BetheLog3S}
\ln k_0(3{S}) &=& 2.767~663~612~491~821~190(1)\,,\\[1ex]
\label{BetheLog4S}
\ln k_0(4{S}) &=& 2.749~811~840~454~057~422(1)\,.
\end{eqnarray}
\end{subequations}
The results for $1S$, $2S$ and $3S$ are in agreement with values 
indicated for the ``logarithmic sum'' $\beta_1$ in~\cite[Tab.~III]{GoDr2000}.
Notice that the numerical difference of our above results
for $\ln k_0$ in comparison to the 
results indicated for $\beta_1$ in~\cite{GoDr2000} is entirely due to an 
additional contribution $\ln 2$ which is added to the 
Bethe logarithm of $S$ states according to the 
somewhat non-standard convention for $\beta_1$ used in~\cite{GoDr2000}.

\begin{table}[htb]
\begin{center}
\begin{minipage}{8.6cm}
\begin{center}
\caption{\label{tableF4S} Numerical results for
the scaled self-energy function
$F(4S_{1/2}, Z\alpha)$ and the self-energy remainder function $G_{\rm SE}$.}
\begin{tabular}{l@{\hspace*{0.5cm}}r@{.}l@{\hspace*{0.5cm}}r@{.}l}
\hline
\hline
$Z$ &
\multicolumn{2}{l}{\rule[-3mm]{0mm}{8mm}
  $F(4{S}_{1/2},Z\alpha)$} &
\multicolumn{2}{l}{\rule[-3mm]{0mm}{8mm}
  $G_{\rm SE}(4{S}_{1/2},Z\alpha)$}\\
\hline
\rule[-1mm]{0mm}{6mm}
$1$ & $10$ & $629~388~4(2)$ & $-30$ & $912(4)$\\
\rule[-1mm]{0mm}{6mm}
$2$ &  $8$ & $841~324~1(2)$ & $-30$ & $380~0(9)$ \\
\rule[-1mm]{0mm}{6mm}
$3$ &  $7$ & $818~078~5(2)$ & $-29$ & $892~4(4)$ \\
\rule[-1mm]{0mm}{6mm}
$4$ &  $7$ & $107~116~6(2)$ & $-29$ & $437~1(2)$ \\
\rule[-3mm]{0mm}{8mm}
$5$ &  $6$ & $566~758~8(2)$ & $-29$ & $006~0(2)$ \\
\hline
\hline
\end{tabular}
\end{center}
\end{minipage}
\end{center}
\end{table}

The evaluation of the coefficient 
\begin{equation}
A_{60}(nS_{1/2}) \equiv \lim_{Z\alpha \to 0} G_{\rm SE}(nS_{1/2},Z\alpha)
\end{equation}
has been
historically problematic~\cite{ErYe1965a,ErYe1965b,%,
Er1971,Sa1981,Pa1993}, and it has therefore been 
a considerable challenge to reliably estimate the 
self-energy remainder function $G_{\rm SE}$,
especially in the 
range of low nuclear charge number $Z$.
Our calculation of the nonperturbative (in 
$Z\alpha$) electron self-energy for the $3{S}_{1/2}$ state 
(see Table~\ref{tableF3S}) has 
a numerical uncertainty of 2~Hz in atomic
hydrogen. For the $4{S}_{1/2}$ state, the
numerical uncertainty is 3$\times Z^4$~Hz (see Table~\ref{tableF4S}).
The value of the fine-structure constant 
$\alpha$ employed in the calculation is
$\alpha^{-1} = 137.036$; this is close to the 
1998 and 2002 CODATA recommended values~\cite{MoTa2000,MoTaPriv2004}. 
The entries for the self-energy remainder 
function $G_{\rm SE}$ in Tables~\ref{tableF3S}
and~\ref{tableF4S} are in agreement with those used 
in the latest adjustment of the fundamental physical 
constants~\cite{MoTaPriv2004} (the $G_{\rm SE}$-values 
used in~\cite{MoTaPriv2004} are based on 
an extrapolation of numerical data previously
obtained~\cite{MoKi1992} for higher nuclear charge numbers). Our all-order
evaluation eliminates any uncertainty due to the unknown
higher-order analytic terms that contribute to 
the bound electron self-energy of $3S$ and $4S$ 
states [see Eq.~(\ref{defFLOnS})].
This improves our knowledge of the spectrum of
hydrogenlike atoms (e.g.~atomic hydrogen, He$^+$).\\[2ex]

{\bf Acknowledgments.}
U. D. J. thanks the National Institute of Standards and Technology for
kind hospitality during a number of extended research appointments. 
The authors acknowledge E.-O.~LeBigot for help in obtaining numerical
results for selected partial contributions to the electron self-energy,
for the hydrogenic energy levels discussed in this work.

\end{document}